# Who is Responsible? Data, Models, or Regulations, A Comprehensive Survey on Responsible Generative AI for a Sustainable Future

### v6

@@ 4.2.1. Comparative Assessment Protocol
Privacy, Safety & Security NASA (Smith and Kearns, 2020): hazard contribution modes of ML in space applications. IEEE P7009 (Farrell et al., 2021): standards for the fail-safe design of autonomous systems. ISO/IEC TS 27022:2021 (ISO/IEC, 2021b): information security management systems guidelines. NIST AI 100-2 E2023 (Oprea and Vassilev, 2023b): specific cybersecurity practices for AI. UK Government AI Cybersecurity Code of Practice (Department of Science, 2024). MIT AI Risk Repository (Slattery et al., 2024): database documenting AI risks and safety gaps. Google’s Secure AI Framework (SAIF) (Google, 2023). Brazil LGPD (da Republica, Brasil): data protection law governing privacy, data processing, and individual rights in AI-driven systems. Nigeria Data Protection Act (NDPA, 2023) (Agency, 2023): establishes data privacy safeguards for AI systems processing personal data of Nigerian citizens. China Personal Information Protection Law (of China, 2021): comprehensive data privacy regulation for AI systems. UAE Federal Decree-Law No. 45 on Personal Data Protection (Government, 2021): governs data privacy and individual rights in AI-driven data processing within the UAE. Unitary (Ltd., [n. d.]): context-aware toxicity tracker. Adversarial Robustness Toolbox (ART) (Research, [n. d.]). Privacy Meter (Privacy and Lab, [n. d.]): audit privacy of ML models. Google Differential Privacy Libraries (Google, [n. d.]). SecretFlow (Group, [n. d.]): privacy-preserving data analysis and ML toolkit. TensorFlow Privacy (Google, [n. d.]). Betterdata.AI (inc., 2024): PET on synthetic data generation.
@@
Ethical Guidelines & Compliance European General Data Protection Regulation (GDPR) (European Parliament and Council of the European Union, 2016). Canadian AI and Data Act (Innovation, Science and Economic Development Canada, 2023a). Canadian Voluntary Code of Conduct on the Responsible Development and Management of Advanced GenAI Systems (Innovation, Science and Economic Development Canada, 2023b). American Executive Order on Safe, Secure, and Trustworthy AI (House, 2023). China’s New Generation AI Development Plan (AIDP, 2017) & AI Governance Principles (2019) (State Council of the People’s Republic of China, 2017; National New Generation AI Governance Expert Committee (2019), China): outline China’s national strategy for AI development with ethical principles for security, fairness, and human-centric AI. UAE National AI Strategy 2031 (United Arab Emirates Government, [n. d.]): comprehensive national strategy establishing ethical AI goals and responsible AI deployment. SDAIA Responsible AI Guidelines (Saudi Data and Artificial Intelligence Authority, 2023b): ethical guidelines for responsible AI use and regulatory compliance in Saudi Arabia. Singapore Model AI Governance Framework (2nd Ed., IMDA) (Infocomm Media Development Authority and Personal Data Protection Commission (PDPC)(2020), IMDA): practical guidance on responsible AI deployment covering internal governance and human oversight. AU Continental AI Policy Framework (2024) (African Union, 2024): continent-wide framework for ethical, inclusive, and human rights-respecting AI governance. Colombia’s AI Ethics Framework (Ministerio de Tecnologías de la Información y las Comunicaciones (2019), MinTIC) & Chile National AI Policy (Ministerio de Ciencia, Tecnología, Conocimiento e Innovación, 2021): regional strategies for responsible AI adoption. Google Responsible AI Practices (Google, [n. d.]). Australia’s AI Ethics Principles (Australian Government Department of Industry, Science and Resources, [n. d.]). Microsoft’s Responsible AI Toolbox (principles and approach, 2024). OneTrust (OneTrust LLC, [n. d.]): demonstrate trust, measure and manage risk, and compliance. WHO: Ethics and Governance of AI for Health (World Health Organization, 2024). Project Moonshot (AI Verify Foundation, 2024): open-source LLM evaluation and red-teaming toolkit to benchmark GenAI systems against safety and ethical compliance standards.
@@
Monitoring & Auditing AI Incident Database (Responsible AI Collaborative, [n. d.]): index of harms realized in the real world by AI systems. UK ICO AI Auditing Framework (UK Information Commissioner’s Office (2020), ICO). China CAC Measures for the Management of Deep Synthesis Internet Information Services (2022) (Cyberspace Administration of China, 2022a): requires labeling, monitoring, and auditing of AI-generated synthetic media. Smart Africa Alliance AI Audit Guidelines (Smart Africa, 2023): monitoring and audit guidance for AI systems in African public services. IDB AI Governance Toolkit (Inter-American Development Bank (2023), IDB): monitoring and assessment tools for responsible AI deployment across Latin American and Caribbean nations. FairVis (Poloclub, [n. d.]): audit classification models for intersectional bias. Aequitas (Center for Data Science and Public Policy, [n. d.]): bias and fairness audit toolkit. Audit-AI : auditing of bias in generalized ML applications.
@@
Table 6. Governance alignment crosswalk (illustrative). ✓= supported evidence; $\triangle$= partial; –= not covered. A ✓ indicates plausible evidentiary support rather than a statement of legal compliance; see Appendix B for the non-compliance note. This crosswalk expands on rubric criteria C9 (metric validity/documentation) and C10 (governance alignment) from Table 4.
@@
NIST: Privacy & Security (of Standards and, NIST, NIST) $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$
@@
Adaptive/Red-team Evidence (of Standards and, NIST) $\triangle$ $\triangle$ $\triangle$ $\triangle$
Abbreviations RTP = RealToxicityPrompts; MM–Saf. = MM–SafetyBench; HarmB. = HarmBench; MLC v0.5 = MLCommons AI Safety v0.5; Humani = HumaniBench.
@@
Building on the challenges in Section 3.2, we apply the rubric in Section 4.2.1 to evaluate benchmarks against governance needs. Table 3 presents the risk-surface matrix for criteria C1–C8. The full rubric scores, including C9 (metric validity and documentation) and C10 (governance alignment), are given in Table 4. Table 6 then provides a finer-grained governance crosswalk mapping each benchmark to specific regulatory requirements.
Risk-surface coverage.
@@
Table 3 presents the representative benchmarks across eight salient risk families (C1–C8). Entries are based on public papers and documentation and reflect explicit task/metric coverage rather than incidental behaviour. The results in Table 3 show that coverage is densest for toxicity and bias (e.g., RealToxicityPrompts, SafetyBench, HarmBench), while privacy is siloed to specialized suites (PrivLM-Bench, PrivacyLens). Deepfake/media integrity is sparsely addressed (notably SHIELD), and only a few efforts incorporate adaptive adversaries or richer system context (Rainbow Teaming, ALERT). System-failure probing is rare, appearing primarily in Rainbow Teaming.
Rubric-scored comparison.
@@
We operationalize the ten-criterion rubric in Section 4.2.1 to produce a normalized $[0,1]$ score per benchmark. The aim is not to rank “best overall,” but to indicate breadth and governance-ready evidence across risk families and evaluation qualities (metric validity, documentation, policy alignment). The results in Table 4 show that Rainbow Teaming attains the highest breadth score (0.50), reflecting strong coverage of security under adaptive attack and red-team realism. Risk Taxonomy and HumaniBench (0.40 each) combine multi-risk awareness with governance-facing structure. MLCommons AI Safety v0.5 and DecodingTrust (0.35) provide broad but mostly static probes. Single-issue suites score lower by design, e.g., RealToxicityPrompts and PrivacyLens (0.15), useful for their focus but limited for governance evidence.
Governance context and crosswalk.
@@
Table 5 situates the analysis within widely used frameworks and tools. Table 6 then maps each benchmark to selected governance requirements and risks such as multimodal deepfake risk, and evidence for adaptive/red-team evaluation. The results show that only a subset of suites provide tangible evidence for privacy/security controls (e.g., PrivLM-Bench). Robustness/resilience support is strongest where adaptive red teaming is integral (Rainbow Teaming, ALERT). Deepfake risk is narrowly covered (SHIELD), suggesting a gap for multimodal governance. Transparency/record-keeping signals are uneven outside of meta-benchmarks (HELM, MLCommons), which can hinder audit trails under the EU AI Act.
@@
Figure 8. Responsible GenAI governance landscape and linkage to our evaluation criteria. Technical factors correspond primarily to rubric items C1–C7 (fairness/bias, toxicity, security/robustness, mis/disinformation, deepfakes/media integrity, privacy, system-level failure). Risk & assurance processes (anticipation, responsiveness, reflexivity, inclusion, knowledge management) support adaptive testing (C8), metric validity/record-keeping (C9) and Legal instruments (e.g., EU AI Act, NIST AI RMF) ground C10 (governance alignment). Sustainable AI highlights environmental/societal/economic impacts that are only weakly covered by current benchmarks, motivating the gaps observed in Tables 46
@@ 4.2.3. Proposed Improvements to AI Safety Benchmarks
Based on the benchmark gaps we found in the literature, we suggest constructive opportunities to complement existing benchmarks and to support broader governance use. First,we recommend that benchmark developers encourage diverse and well-documented corpora across regions, languages, and demographics, accompanied by dataset cards or data statements that capture provenance, consent, licensing, and curation criteria to enhance traceability and accountability (Eriksson et al., 2025; Gebru et al., 2021). Equally important is construct clarity: each task and metric should state explicitly what it is intended to measure and, where feasible, be validated with multidisciplinary input to guard against superficial compliance claims (Ren et al., 2024). Without these foundations, high benchmark scores risk measuring test-taking ability rather than genuine safety.
Incorporating adversarial evaluation, including undisclosed items and evolving prompts can be used bring benchmarks in line with established security testing practices (Oprea and Vassilev, 2023b; Ghosh et al., 2025). Scope should also expand to cover multimodal, non-English, and higher-risk domains, as well as multi-agent or autonomous settings where system-level failures are more likely to surface (Salazar et al., 2025). These benchmarks rarely make explicit linkages to policy instruments. Engaging multiple stakeholders and mapping evaluation outputs to governance frameworks would mitigate commercial or domain bias and improve regulatory utility (Outeda, 2024). Continuous evaluation through automated pipelines that enable periodic re-testing across the model lifecycle would further improve comparability over time (Wang et al., 2024). Finally, structured red teaming and stress tests under distribution shift should be integrated to surface emergent behaviours that may elude static suites, drawing on practices such as ALERT and Rainbow Teaming.
@@ 4.2.4. Key Performance Indicators (KPIs) for Responsible GenAI
KPIs provide measurable criteria to assess the performance of AI systems across critical dimensions of responsibility (Mattioli et al., 2024). Unlike ethical principles that are mostly subjective, KPIs translate RAI goals into operational metrics that can be tracked, compared, and audited across the model lifecycle. They bridge governance expectations with technical implementation so that fairness, transparency, robustness, privacy, and sustainability are evidenced through quantifiable measures, enabling compliance and continuous improvement. We adapt established measures (SPD, DI, accuracy parity, robustness under perturbation, and energy) from prior work (Pessach and Shmueli, 2022), and introduce pragmatic, author-defined KPIs for explainability, inclusivity, auditing cadence, and high-stakes errors to support governance-ready evidence.
@@
    $Q=\frac{\mathrm{Valid\ Entries}}{\mathrm{Total\ Entries}}\times 100.$
@@
    $P=\frac{\mathrm{Compliant\ Data\ Points}}{\mathrm{Total\ Data\ Points}}\times 100.$
@@
    $\Delta_{\mathrm{acc}}=\left|\mathrm{Accuracy}_{A=a}-\mathrm{Accuracy}_{A=b}\right|.$
@@
    $R=\frac{\mathrm{Errors\ under\ perturbation}}{\mathrm{Total\ perturbed\ inputs}}\times 100,\quad\Delta_{\mathrm{acc}}^{\mathrm{rob}}=\mathrm{Acc}_{\mathrm{clean}}-\mathrm{Acc}_{\mathrm{pert}}.$
@@
    $\mathrm{LFG}(x)=\frac{\left|\,f(x)-\left(\phi_{0}+\sum_{i}\phi_{i}(x)\right)\right|}{\max\left(\left|f(x)\right|,\epsilon\right)}.$
@@
    $E_{h}=\frac{\mathrm{Critical\ errors}}{\mathrm{Total\ predictions}}\times 100.$
@@
    $F_{a}=\frac{\mathrm{Audits}}{\mathrm{Time\ period}},\qquad T_{r}=\mathrm{median\ time\ to\ resolve}.$
@@
    $U=\frac{\mathrm{Inclusive\ features\ implemented}}{\mathrm{Planned\ features}}\times 100.$
@@
    $E_{\mathrm{kWh}}=\frac{P_{\mathrm{avg}}}{1000}\times T,\qquad C_{\mathrm{CO_{2}e}}=E_{\mathrm{kWh}}\times\mathrm{PUE}\times\mathrm{CI}.$

### v7

@@ 4.2.1. Comparative Assessment Protocol
Privacy, Safety & Security NASA (Smith and Kearns, 2020): hazard contribution modes of ML in space applications. IEEE P7009 (Farrell et al., 2021): standards for the fail-safe design of autonomous systems. ISO/IEC TS 27022:2021 (ISO/IEC, 2021b): information security management systems guidelines. NIST AI 100-2 E2023 (Oprea and Vassilev, 2023): specific cybersecurity practices for AI. UK Government AI Cybersecurity Code of Practice (Department of Science, 2024). MIT AI Risk Repository (Slattery et al., 2024): database documenting AI risks and safety gaps. Google’s Secure AI Framework (SAIF) (Google, 2023). Brazil LGPD (da Republica (Brasil), 2018): data protection law governing privacy, data processing, and individual rights in AI-driven systems. Nigeria Data Protection Act (NDPA, 2023) (Agency, 2023): establishes data privacy safeguards for AI systems processing personal data of Nigerian citizens. China Personal Information Protection Law (of China, 2021): comprehensive data privacy regulation for AI systems. UAE Federal Decree-Law No. 45 on Personal Data Protection (UAE Government, 2021): governs data privacy and individual rights in AI-driven data processing within the UAE. Unitary (Ltd., ): context-aware toxicity tracker. Adversarial Robustness Toolbox (ART) (Research, ). Privacy Meter (Privacy and Lab, ): audit privacy of ML models. SecretFlow (Group, ): privacy-preserving data analysis and ML toolkit. Betterdata.AI (inc., 2024): PET on synthetic data generation.
@@
Ethical Guidelines & Compliance European General Data Protection Regulation (GDPR) (European Parliament and Council of the European Union, 2016). Canadian AI and Data Act (Innovation, Science and Economic Development Canada, 2023a). Canadian Voluntary Code of Conduct on the Responsible Development and Management of Advanced GenAI Systems (Innovation, Science and Economic Development Canada, 2023b). American Executive Order on Safe, Secure, and Trustworthy AI (House, 2023). China’s New Generation AI Development Plan (AIDP, 2017) & AI Governance Principles (2019) (State Council of the People’s Republic of China, 2017; National New Generation AI Governance Expert Committee (China), 2019): outline China’s national strategy for AI development with ethical principles for security, fairness, and human-centric AI. UAE National AI Strategy 2031 (UAE Government, ): comprehensive national strategy establishing ethical AI goals and responsible AI deployment. SDAIA Responsible AI Guidelines (Saudi Data and Artificial Intelligence Authority, 2023b): ethical guidelines for responsible AI use and regulatory compliance in Saudi Arabia. Singapore Model AI Governance Framework (2nd Ed., IMDA) (Infocomm Media Development Authority (IMDA) and Personal Data Protection Commission (PDPC), 2020): practical guidance on responsible AI deployment covering internal governance and human oversight. Colombia’s AI Ethics Framework (Ministerio de Tecnologías de la Información y las Comunicaciones (MinTIC), 2019) & Chile National AI Policy (Ministerio de Ciencia, Tecnología, Conocimiento e Innovación, 2021): regional strategies for responsible AI adoption. Google Responsible AI Practices (Google, ). Australia’s AI Ethics Principles (Australian Government Department of Industry, Science and Resources, ). Microsoft’s Responsible AI Toolbox (principles and approach, 2024). OneTrust (OneTrust LLC, ): demonstrate trust, measure and manage risk, and compliance. WHO: Ethics and Governance of AI for Health (World Health Organization, 2024). Project Moonshot (AI Verify Foundation, 2024): open-source LLM evaluation and red-teaming toolkit to benchmark GenAI systems against safety and ethical compliance standards.
@@
Monitoring & Auditing AI Incident Database (Responsible AI Collaborative, ): index of harms realized in the real world by AI systems. UK ICO AI Auditing Framework (UK Information Commissioner’s Office (ICO), 2020). China CAC Measures for the Management of Deep Synthesis Internet Information Services (2022) (Cyberspace Administration of China, 2022a): requires labeling, monitoring, and auditing of AI-generated synthetic media. Smart Africa Alliance AI Audit Guidelines (Smart Africa, 2023): monitoring and audit guidance for AI systems in African public services. IDB AI Governance Toolkit (Inter-American Development Bank (IDB), 2023): monitoring and assessment tools for responsible AI deployment across Latin American and Caribbean nations. FairVis (Poloclub, ): audit classification models for intersectional bias. Aequitas (Center for Data Science and Public Policy, ): bias and fairness audit toolkit. Audit-AI : auditing of bias in generalized ML applications.
@@
Table 5. Governance alignment crosswalk (illustrative). ✓= supported evidence; $\triangle$= partial; –= not covered. A ✓ indicates plausible evidentiary support rather than a statement of legal compliance; see Appendix 1.3 for the non-compliance note. This crosswalk expands on rubric criteria C9 (metric validity/documentation) and C10 (governance alignment) from Table 3.
@@
NIST: Privacy & Security (of Standards and (NIST), 2020a, b) $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$
@@
Adaptive/Red-team Evidence (of Standards and (NIST), 2023) $\triangle$ $\triangle$ $\triangle$ $\triangle$
RTP = RealToxicityPrompts; MM–Saf. = MM–SafetyBench; HarmB. = HarmBench; MLC v0.5 = MLCommons AI Safety v0.5; Humani = HumaniBench.
@@
Building on the challenges in Section 3.2, we apply the rubric in Section 4.2.1 to evaluate benchmarks against governance needs and discuss our findings below.
@@
Risk-surface coverage: Table 3 presents the representative benchmarks across eight salient risk families (C1–C8). Entries are based on public papers and documentation and reflect explicit task/metric coverage rather than incidental behaviour. Table 3 alsoshow that coverage is densest for toxicity and bias (e.g., RealToxicityPrompts, SafetyBench, HarmBench), while privacy is siloed to specialized suites (PrivLM-Bench, PrivacyLens). Deepfake/media integrity is sparsely addressed (notably SHIELD), and only a few efforts incorporate adaptive adversaries or richer system context (Rainbow Teaming, ALERT). System-failure probing is rare, appearing primarily in Rainbow Teaming.
@@
Rubric-scored comparison: We operationalize the ten-criterion rubric in Section 4.2.1 to produce a normalized $[0,1]$ score per benchmark. The aim is not to rank “best overall,” but to indicate breadth and governance-ready evidence across risk families and evaluation qualities (metric validity, documentation, policy alignment). The results in Table 3 show that Rainbow Teaming attains the highest breadth score (0.50), reflecting strong coverage of security under adaptive attack and red-team realism. Risk Taxonomy and HumaniBench (0.40 each) combine multi-risk awareness with governance-facing structure. MLCommons AI Safety v0.5 and DecodingTrust (0.35) provide broad but mostly static probes. Single-issue suites score lower by design, e.g., RealToxicityPrompts and PrivacyLens (0.15), useful for their focus but limited for governance evidence.
@@
Governance context and crosswalk: Table 4 situates the analysis within widely used frameworks and tools. Table 5 then maps each benchmark to selected governance requirements and risks such as multimodal deepfake risk, and evidence for adaptive/red-team evaluation. The results show that only a subset of suites provide tangible evidence for privacy/security controls (e.g., PrivLM-Bench). Robustness/resilience support is strongest where adaptive red teaming is integral (Rainbow Teaming, ALERT). Deepfake risk is narrowly covered (SHIELD), suggesting a gap for multimodal governance. Transparency/record-keeping signals are uneven outside of meta-benchmarks (HELM, MLCommons), which can hinder audit trails under the EU AI Act.
@@
Figure 7. Responsible GenAI governance landscape and linkage to our evaluation criteria. Technical factors correspond primarily to rubric items C1–C7 (fairness/bias, toxicity, security/robustness, mis/disinformation, deepfakes/media integrity, privacy, system-level failure). Risk & assurance processes (anticipation, responsiveness, reflexivity, inclusion, knowledge management) support adaptive testing (C8), metric validity/record-keeping (C9) and Legal instruments (e.g., EU AI Act, NIST AI RMF) ground C10 (governance alignment). Sustainable AI highlights environmental/societal/economic impacts that are only weakly covered by current benchmarks.
@@ 4.2.3. Proposed Improvements to AI Safety Benchmarks
Based on the benchmark gaps we found in the literature, we suggest constructive opportunities to complement existing benchmarks and to support broader governance use. First, we recommend that benchmark developers encourage diverse and well-documented corpora across regions, languages, and demographics, accompanied by dataset cards or data statements that capture provenance, consent, licensing, and curation criteria to enhance traceability and accountability (Gebru et al., 2021). Also, each task and metric should state explicitly what it is intended to measure and, where feasible, be validated with multidisciplinary input to guard against superficial compliance claims (Ren et al., 2024).
Incorporating adversarial evaluation, including undisclosed items and evolving prompts can be used bring benchmarks in line with established security testing practices. Scope should also expand to cover multimodal, non-English, and higher-risk domains, as well as multi-agent or autonomous settings where system-level failures are more likely to surface. Engaging multiple stakeholders and mapping evaluation outputs to governance frameworks can help mitigate commercial or domain bias and improve regulatory utility. Finally, structured red teaming and stress tests under distribution shift should be integrated to surface emergent behaviours that may elude static suites, drawing on practices such as ALERT and Rainbow Teaming.
@@ 4.2.4. Key Performance Indicators (KPIs) for Responsible GenAI
KPIs provide measurable criteria to assess the performance of AI systems across critical dimensions of responsibility (Mattioli et al., 2024). Unlike ethical principles that are mostly subjective, KPIs translate RAI goals into operational metrics that can be tracked, compared, and audited across the model lifecycle. They bridge governance expectations with technical implementation so that fairness, transparency, robustness, privacy, and sustainability are evidenced through quantifiable measures, enabling compliance and continuous improvement. We adapt established measures (SPD, DI, accuracy parity, robustness under perturbation, and energy) from prior works (Pessach and Shmueli, 2022; Raza et al., 2024), and introduce KPIs for explainability, inclusivity, auditing cadence, and high-stakes errors to support governance-ready evidence.
@@
    (2) $Q=\frac{\mathrm{Valid\ Entries}}{\mathrm{Total\ Entries}}\times 100.$
@@
    (3) $P=\frac{\mathrm{Compliant\ Data\ Points}}{\mathrm{Total\ Data\ Points}}\times 100.$
@@
    (6) $\Delta_{\mathrm{acc}}=\left|\mathrm{Accuracy}_{A=a}-\mathrm{Accuracy}_{A=b}\right|.$
@@
    (7) $R=\frac{\mathrm{Errors\ under\ perturbation}}{\mathrm{Total\ perturbed\ inputs}}\times 100,\quad\Delta_{\mathrm{acc}}^{\mathrm{rob}}=\mathrm{Acc}_{\mathrm{clean}}-\mathrm{Acc}_{\mathrm{pert}}.$
@@
    (9) $\mathrm{LFG}(x)=\frac{\left|\,f(x)-\left(\phi_{0}+\sum_{i}\phi_{i}(x)\right)\right|}{\max\left(\left|f(x)\right|,\epsilon\right)}.$
@@
    (10) $E_{h}=\frac{\mathrm{Critical\ errors}}{\mathrm{Total\ predictions}}\times 100.$
@@
    (11) $F_{a}=\frac{\mathrm{Audits}}{\mathrm{Time\ period}},\qquad T_{r}=\mathrm{median\ time\ to\ resolve}.$
@@
    (12) $U=\frac{\mathrm{Inclusive\ features\ implemented}}{\mathrm{Planned\ features}}\times 100.$
@@
    (13) $E_{\mathrm{kWh}}=\frac{P_{\mathrm{avg}}}{1000}\times T,\qquad C_{\mathrm{CO_{2}e}}=E_{\mathrm{kWh}}\times\mathrm{PUE}\times\mathrm{CI}.$
@@
Recent GenAI era deployments expose failure modes (hallucination, privacy leakage, deepfakes, and tool-use errors) that cut across domains; we therefore use C1–C10 and the KPI set as a consistent lens throughout the remainder of the survey.